# Learning System Dynamics without Forgetting

### v1

@@ 4.5 Comparisons with State-of-the-Arts (RQ3,4)
In this subsection, we compare MS-GODE with multiple state-of-the-arts methods including the joint training, which is typically regarded as the upper bound on the performance in continual learning research. The experiments are conducted on both physics systems 1 and cellular systems 2, which demonstrate that MS-GODE outperforms the baselines in both cellular systems and physics systems with different configurations. Besides, by comparing the results across different sequences in the two tables, we could find that the sequences with gradual dynamics shift (Sequence 1,2 of physics systems) are more difficult to learn than the sequences with abrupt dynamics shift (Sequence 3 of the physics systems and the cellular system).
@@ 4.6 In-depth Investigation on the Learning Dynamics (RQ3,4)
Table 1 and 2 provide the overall performance, which is convenient to compare different methods. However, as introduced in Section 4.2, to obtain an in-depth understanding of the performance of different methods, we have to seek help from the most through metric, i.e. the performance matrix. In Figure 6, we visualize the performance marices of different methods after learning different system sequences. The $i$-th column demonstrates the performance of the $i$-th task when learning sequentially over the systems. Comparing MS-GODE ((a) and (e)) and the other methods, we find that MS-GODE could maintain a much more stable performance of each system when learning over the sequence. EWC ((c) and (f)) also maintains a relatively stable performance of each system based on its regularization strategy. However, compared to MS-GODE, the model becomes less and less adaptive to new systems (the columns become increasingly darker from left to right). This is because the regularization is applied to more parameters when proceeding to each new task. Fine-tune ((b) and (g)) is not limited by regularization, therefore is more adaptive on new systems but less capable in preserving the performance on previous systems compared to EWC. LwF (d), although based on knowledge distillation, does not directly limit the adaptation of the parameters like EWC. Finally, based on memory and gradient modification, GEM (h) maintains the performance better than fine-tune (g), and is more adaptive to new tasks than EWC (f). More details on the performance matrices is provided in Appendix A.5.
Figure 6: Visualization of performance matrices. (a), (b), (c), (d) are performance of MS-GODE, fine-tune, EWC, and LwF on the cellular system sequence. (e), (f), (g), (h) are the performance of MS-GODE, EWC, fine-tune, and GEM, on the physics systems.
@@ 5 Conclusion
In this paper, we systematically study the problem of continually learning varying system dynamics, and propose an effective method MS-GODE for this task. Moreover, we also construct a novel benchmark consisting of biological cellular systems for the field of machine learning on system dynamics. Finally, we conduct comprehensive experiments on both physics and cellular system sequences, which not only demonstrate the effectiveness of MS-GODE, but also provide insights into the problem of machine learning over system sequences with dynamics shift. Although MS-GODE outperforms the baselines significantly, there is still a gap between MS-GODE and the performance upper bound (joint training) especially in the biological cellular systems. In our future work, we aim to develop improved strategies for sub-network learning to further boost the performance in both individual systems and system sequences.
@@ A.1.1 Physics System Sequence
Simulated physics systems are commonly adopted to evaluate the machine learning models in the task of learning system dynamics [6, 2, 7]. The physics systems adopted in this work include spring connected particles and charged particles (Figure 2) with disparate dynamics, therefore are ideal for constructing system sequences to evaluate a model’s continual learning capability under severe significant dynamics shift. Besides, the configuration of each system type is also adjustable. For the spring connected particles, the number of particles, strength of the springs, and the size of the box containing the particles are adjustable. For the charged particles, the number of particles, charge sign, and the size of box are adjustable. In our experiments, we constructed multiple systems with different configurations, which are aligned into sequences for the model to learn.
@@ A.1.2 Biological Cellular System
In this paper, we build a novel benchmark containing biological cellular dynamic systems based on Virtual Cell [13, 14, 15] with different system configurations and variable selection. Currently, the benchmark is built based on two types of cellular models. The first one is rule-based model of EGFR receptor interaction with two adapter proteins Grb2 and Shc. The second is a compartmental rule based model of translocation through the nuclear pore of a cargo protein based on the Ran protein (GTPase). Detailed description of these two systems can be found via https://vcell.org/webstart/VCell_Tutorials/VCell6.1_Rule-Based_Tutorial.pdf and https://vcell.org/webstart/VCell_Tutorials/VCell6.1_Rule-Based_Ran_Transport_Tutorial.pdf. Based on these 2 types of models, we construct multiple systems, which are aligned into different system sequences.
@@
In our experiments, we adjust the parameter configurations of these two types of models to construct multiple systems with different dynamics, which are aligned into a sequence for the experiments. We generate simulated data using the Virtual Cell platform (VCell) and pre-process the data. In experiments, we adopt a 4-system sequence using the simulated data by alternatively align 2 EGFR systems and 2 Ran systems. EGFR system is rule-based model of EGFR receptor interaction with two adapter proteins Grb2 and Shc. Ran system is a compartmental rule based model of translocation through the nuclear pore of a cargo protein based on the Ran protein (GTPase).
@@
For each cellular system, we generate 100 iterations of simulation using VCell. Different from the physics systems, the temporal interval between two time steps of cellular system is not constant. Similar to the physics systems, the first 60% part of the observation of each system simulation is fed into the model to reconstruct the remaining 40% during training. For each system, we generate 20 systems for training and 20 systems for testing.
@@ A.1.3 Additional Details of Experimental Settings
For the encoder network, the number of layers is 2, and the hidden dimension is 64. 1 head is used for the attention mechanism. The interaction network of the generator is configured as 1-layer network, and the number of hidden dimensions is 128. Finally, the decoder network is a fully-connected layer. The model is trained for 20 epochs over each system in the given sequence. We adopt the AdamW optimizer [47] and set the learning rate as 0.0005.
@@ A.2 Edge-popup Algorithm
In this work, we adopt the edge-popup algorithm [11] for optimizing the binary masks. The main idea is to optimize a continuous score value for each entry of the mask during the backpropagation, and binarize the values into discrete binary values during forward propagation (Appendix A.2). Accordingly, the strategy for binarizing the mask entry values is a crucial factor influencing the performance. For convenience of the readers, we provide the details about this algorithm in this subsection.
@@
where $h(s_{uv})$ is the binary value over the weight $w_{uv}$ denoting whether this weight is selected and $s_{uv}$ the continuous score used in gradient descent based optimization. During backpropagatioon, the gradient will ignore $h(\cdot)$ and goes through it, therefore the gradient over $s_{uv}$ is
@@ A.3 Baselines
    Elastic Weight Consolidation (EWC) [28] applies a quadratic penalty over the parameters of a model based on their importance to maintain its performance on previous tasks.
@@
    Gradient Episodic Memory (GEM) [34] selects and stores representative data in an episodic memory buffer. During training, GEM will modify the gradients calculated based on the current task with the gradient calculated based on the stored data to avoid updating the model into a direction that is detrimental to the performance on previous tasks.
@@
    Learning without Forgetting (LwF) [44] is a knowledge distillation based method, which minimizes the discrepancy between the the old model output and the new model output to preserve the knowledge learned from the old tasks.
@@
    Bias Correction based Memory Replay (BCMR) [45]. This baseline is constructed by integrating the navie memory replay with the data sampling bias correction strategy  [45]. In other words, the method does not train the data immediately after observing the data. Instead, it stores the observed data into a memory buffer. Whenever testing is required, the model will be trained over all buffered data.
@@
    Scheduled Data Prior (SDP) [46] considers that the importance of new and old data is dependent on the specific characteristics of the given data, therefore balance the contribution of new and old data based on a data-driven approach.
@@ A.4 Model Evaluation
Different from standard learning setting with only one task to learn and evaluate, in our setting, the model will continually learn on a sequence of systems, therefore the setting and evaluation are significantly different. In the model training stage, the model is trained over a system sequence. In the testing stage, the model will be tested on all learned tasks. Therefore the model will have multiple performance corresponding to different tasks, and the most thorough evaluation metric is the performance matrix $\mathrm{M}^{p}\in\mathbb{R}^{N\times N}$, where $\mathrm{M}^{p}_{i,j}$ denotes the performance on the $j$-system after learning from the $1$-st to the $i$-th system, and $N$ is the number of systems in the sequence. In our experiments, each entry of $\mathrm{M}^{p}$ is a mean square error (MSE) value. To evaluate the overall performance on a sequence, the average performance (AM) over all learnt tasks after learning multiple tasks could be calculated. For example, i.e., $\frac{\sum_{j=1}^{i}\mathrm{M}^{p}_{N,j}}{N}$ corresponds to the average model performance after learning the entire sequence with $N$ tasks. Similarly, the average forgetting (AF) after $N$ tasks can be formulated as $\frac{\sum_{j=1}^{N-1}\mathrm{M}^{p}_{N,j}-\mathrm{M}^{p}_{j,j}}{N-1}$. These metrics are widely adopted in continual learning works [48, 34, 49, 50, 51], although the names are different in different works.
For convenience, the performance matrix can be visualized as a color map (Figure 6). For example, they are named as Average Accuracy (ACC) and Backwarde Transfer (BWT) in [48, 34], Average Performance (AP) and Average Forgetting (AF) in [49], Accuracy Mean (AM) and Forgetting Mean (FM) in [50], and performance mean (PM) and forgetting mean (FM) in [51].

### v2

@@ 4.5 Comparisons with State-of-the-Arts (RQ3,4)
In this subsection, we compare MS-GODE with state-of-the-arts methods including the joint training. The experiments are conducted on both physics systems (Table 1) and cellular systems (Table 2). Besides, by comparing the results across different sequences in the two tables, we find that the sequences with gradual dynamics shift (Sequence 1,2 of physics systems) are more difficult to learn than the sequences with abrupt dynamics shift (Sequence 3 of the physics systems and the cellular system). The advantages of MS-GODE over the baselines mainly come from two perspectives. First, since the learning of the subnetworks are learned in a completely independent manner, the plasticity when learning new systems is guaranteed. Second, the independency of the subnetworks also eliminates the forgetting issue and protect the performance stability on previously learned systems.
@@ 4.6 In-depth Investigation on the Learning Dynamics (RQ3,4)
Table 1 and 2 provide the overall performance. However, as introduced in Section 4.2, to obtain an in-depth understanding of the performance of different methods, we have to seek help from the performance matrix. In Figure 6, we visualize the performance matrices of different methods. The $i$-th column demonstrates the performance of the $i$-th task when learning sequentially over the systems. Comparing MS-GODE ((a) and (e)) and the other methods, we find that MS-GODE maintains a much more stable performance. EWC ((c) and (f)) also maintains a relatively stable performance of each system based on its regularization strategy. However, compared to MS-GODE, the model becomes less and less adaptive to new systems, because the regularization is applied to more parameters when proceeding to each new task. Fine-tune ((b) and (g)) is not limited by regularization, therefore is more adaptive on new systems but less capable of preserving the performance on previous systems compared to EWC. LwF (d), although based on knowledge distillation, does not directly limit the adaptation of the parameters. Finally, based on memory and gradient modification, GEM (h) maintains the performance better than fine-tune (g), and is more adaptive to new tasks than EWC (f). More details on the performance matrices are provided in Appendix A.5.
Figure 6: Visualization of performance matrices. (a), (b), (c), (d) corresponds to MS-GODE, fine-tune, EWC, and LwF on the cellular system sequence. (e), (f), (g), (h) are the performance of MS-GODE, EWC, fine-tune, and GEM, on the physics systems. Lower values indicate better performance.
@@ 5 Conclusion
In this paper, we systematically study the problem of continual dynamics learning (CDL) from different perspectives, including investigating the influence of task configuration and evaluating the performance of existing continual learning techniques. Based on the findings, we propose Mode-switching Graph ODE (MS-GODE). Additionally, we also construct Bio-CDL, consisting of biological cellular systems and significantly enriching the research field of machine learning on system dynamics. Finally, we conduct comprehensive experiments on both physics and cellular system sequences, which not only demonstrate the effectiveness of MS-GODE, but also provide insights into the problem of machine learning over system sequences with dynamics shift.
Acknowledgments
This project is supported by the National Research Foundation, Singapore, under its NRF Professorship Award No. NRF-P2024-001.
@@ A.1.1 Physics System Sequence
Simulated physics systems are commonly adopted to evaluate the machine learning models in the task of learning system dynamics (Liu et al., 2024; Battaglia et al., 2016; Huang et al., 2020). The physics systems adopted in this work include spring connected particles and charged particles (Figure 7) with disparate dynamics, therefore are ideal for constructing system sequences to evaluate a model’s continual learning capability under severe significant dynamics shift. Besides, the configuration of each system type is also adjustable. For the spring connected particles, the number of particles, strength of the springs, and the size of the box containing the particles are adjustable. For the charged particles, the number of particles, charge sign, and the size of box are adjustable. In our experiments, we constructed multiple systems with different configurations, which are aligned into sequences for the model to learn.
@@ A.1.2 Biological Cellular System
In this paper, we build a novel benchmark, Bio-CDL, containing biological cellular dynamic systems based on Virtual Cell (Schaff et al., 1997; Cowan et al., 2012; Blinov et al., 2017) with different system configurations and variable selection. Currently, Bio-CDL is built based on two types of cellular models. The first one is rule-based model of EGFR receptor interaction with two adapter proteins Grb2 and Shc. The second is a compartmental rule based model of translocation through the nuclear pore of a cargo protein based on the Ran protein (GTPase). An illustration of the Ran system is provided in Figure 8.
Figure 8: Illustration of the RAN-regulated nucleocytoplasmic transport (Moore, 2013). Briefly speaking, this model depicts the translocation of cargo proteins (Exportin 1) via nuclear pores with the assistance of RAN proteins. RAN-GDP is first activated into RAN-GTP) and then binds to cargo molecules (Exportin 1) forming a complex containing RAN-GTP and Exportin 1. Next, the complex is translocated across the nuclear membrane into the cytoplasm with the assistance of RAN. Finally, RAN and Exportin 1 are dissociated after the translocation.
Detailed description of these two systems can be found via https://vcell.org/webstart/VCell_Tutorials/VCell6.1_Rule-Based_Tutorial.pdf and https://vcell.org/webstart/VCell_Tutorials/VCell6.1_Rule-Based_Ran_Transport_Tutorial.pdf. Based on these 2 types of models, we construct multiple systems, which are aligned into different system sequences. Details are provided below.
Our new benchmark is constructed based on the EGFR receptor interaction model and Ran transportation model. The length of the observe trajectories ranges from 250 to 550 time steps for EGFR systems and 40 to 150 time steps for Ran systems.
We provide below the details of data generation, and will release our generated data later. The code for generating the configuration files for the simulations is contained in
    VCell_config_gen.py.
We have restricted the model coefficients to a specific range to ensure the simulations remain stable. After generating the configuration files, they should be uploaded into the VCell platform for generating the simulations. The specific steps are listed below.
[…]
Figure 9:
[…]
Figure 10:
4. In the main window on the upper right, select the ’Simulations’ tab and click the first icon under the ’Simulations’ tab to create a new simulation (Figure 11).
Figure 11:
5. Click to select the created simulation, then click the icon with a blue arrow and a gear to load the previously generated simulations. 6. After the simulations are done, select all the simulations and click the icon with a green arrow and a gear to export all the simulation results into a specified folder. 7. Specify the
     data_store_path
and run
 generate_dataset.py
to obtain the data files that can be loaded and used by the MS-GODE model.
This implementation of MS-GODE is based on [Pytorch Geometric]https://github.com/rusty1s/pytorch_geometric API.
Example command for running the experiments with MS-GODE:
python run_models.py --cut_num 20 \
--nepos 20 \
--device 5 \
--n_iters_to_viz 10 \
--thresholding ’fast’ \
--mask True \
--dropout_mask ’0.0’ \
--batch-size 10 \
--mode "ex" \
--normalizeVCellfeat ’universal’ \
--system ’VCell’ \
--repeats 5 \
--overwrite_results ’False’ \
--fix_random_seed ’False’ \
--save_results True
Example scripts for data generation of cellular/physics systems:
 python ./data/generate_dataset.py --simulation VCell \
  --num-train 3000 \
  --num-test 3000 \
  --n-balls 5
The generation of cellular simulation data requires first obtaining simulation data from VCell platform, which is introduced in the Appendix A.1.2 of the submission.
 python ./data/generate_dataset.py --simulation simulation \
  --num-train 3000 \
  --num-test 3000 \
  --n-balls 5
Setup
The mdoel implementation is based on the following packages:
- [Python 3.6.10](https://www.python.org/)
- [Pytorch 1.4.0](https://pytorch.org/)
[…]
- torch-cluster==1.5.3 - torch-scatter==2.0.4 - torch-sparse==0.6.1
- [torchdiffeq](https://github.com/rtqichen/torchdiffeq)
- [numpy 1.16.1](https://numpy.org/)
@@
In our experiments, we adjust the parameter configurations of these two types of models to construct multiple systems with different dynamics, which are aligned into a sequence for the experiments. We generate simulated data using the Virtual Cell platform (VCell) and pre-process the data. In experiments, we adopt a 4-system sequence using the simulated data by alternatively align 2 EGFR systems and 2 Ran systems. EGFR system is rule-based model of EGFR receptor interaction with two adapter proteins Grb2 and Shc. Ran system is a compartmental rule based model of translocation through the nuclear pore of a cargo protein based on the Ran protein (GTPase). For constructing systems with different dynamics, different EGFR and Ran systems are constructed by selecting a subset of the observed variables from the complete systems. This not only enrich the dynamics diversity across different systems in the learning sequence, but also enhance the learning difficulty over the systems. The IDs (inherited from VCell platform) of the selected variables in $\mathrm{EGFR}_{1}$, $\mathrm{EGFR}_{2}$, $\mathrm{Ran}_{1}$, $\mathrm{Ran}_{2}$ are {s15,s16,s21,s22,s29,s33,s34, s38, s45, s47, s59, s74, s9, s96}, {s24, s25, s26, s27, s30, s36, s43, s44, s58, s68}, {Ran_C_nuc, RCC1, s10, s11, s12, s13, s14, s15, s16, s17, s18, s19, s2, s20, s21, s22, s23, s24, s25, s26, s27, s28, s29, s3, s30, s31, s32, s33, s34, s35, s4, s5, s6, s7, s8, s9 }, {s14, s15, s16, s23, s24, s25, s31, s5 }, respectively.
Using the code for generating cellular system data requires the following steps.
  1. 1.
    Using the script contained in our provided code ‘MS-GODE/VCell_config_gen’ to generate the configuration files for VCell simulation. Users can freely adjust the configurations in the script.
  2. 2.
    Download the VCell application from https://vcell.org/.
  3. 3.
    Load the generated configuration files into the VCell platform.
  4. 4.
    Generate simulation results. Instruction of using this script is also provided in the ReadMe file of our code.
  5. 5.
    Using the script ‘MS-GODE/data/generate_dataset.py’ for processing the simulation results. After this step, the data can be used for experiments.
Since the entire dataset is much larger than the 100 MB limit, we cannot provide the data in the supplementary materials. But we have provided all code and instruction for generating the data, and will release the complete Bio-CDL benchmark to the public later. Besides the sequence $\mathrm{EGFR}_{1}\rightarrow\mathrm{Ran}_{1}\rightarrow\mathrm{EGFR}_{2}\rightarrow\mathrm{Ran}_{2}$ adopted in the experiments, the complete Bio-CDL also contain more different system configurations and sequence configurations. Moreover, we are also working on enriching Bio-CDL with more diverse system types and system sequences.
@@
For each cellular system, we generate 100 iterations of simulation using VCell. Different from the physics systems, the temporal interval between two time steps of cellular system is not constant. Similar to the physics systems, the first 60% part of the observation of each system simulation is fed into the model to reconstruct the remaining 40% during training. For each system, we use 20 trajectories for training and 20 trajectories for testing.
@@ A.1.3 Additional Details of Experimental Settings
For the encoder network, the number of layers is 2, and the hidden dimension is 64. 1 head is used for the attention mechanism. The interaction network of the generator is configured as 1-layer network, and the number of hidden dimensions is 128. Finally, the decoder network is a fully-connected layer. The model is trained for 20 epochs over each system in the given sequence. We adopt the AdamW optimizer (Loshchilov & Hutter, 2017) and set the learning rate as 0.0005. In our work, task boundaries are provided to the models during training.
A.1.4 Additional discussion on performance and low-level dynamics shift
The performance of MS-GODE is mainly determined on two factors, including the performance on each system and the forgetting issue. Therefore, to further improve the performance under gradual dynamics shift, there are two promising approaches:
1. Enhancing the performance on each single system. As revealed in Ramanujan et al. (2020), the larger the backbone network, the more probable the masked sub-network can reach the capacity of the full backbone network. In other words, increasing the size of the model can improve the performance.
2. Reducing the performance decrease after learning new systems. Since MS-GODE completely separate the masks for different dynamics, the forgetting issue has been eliminated. Therefore, the performance decrease after learning new systems mainly comes from the incorrect mask selection. The strategy above to increase model size can help improve the mask selection, because better fitting to each system increases the difference between the masks. Furthermore, the mask selection can also be improved by incorporating a mixture of selection criteria. Specifically, in our experiments, the mode-switching module of MS-GODE selects the mask based on the error of reconstructing the second 50% of the observation. This can be enhanced by incorporating different splitting ratios and mixing the result. Since the correct mask tend to exhibit lower error with most splitting ratios, selecting the one that succeeds in more splitting ratios could increase the possibility of finding the correct one.
@@ A.2 Edge-popup Algorithm
In this work, we adopt the edge-popup algorithm (Ramanujan et al., 2020) for optimizing the binary masks. The main idea is to optimize a continuous score value for each entry of the mask during the backpropagation, and binarize the values into discrete binary values during forward propagation (Appendix A.2). Accordingly, the strategy for binarizing the mask entry values is a crucial factor influencing the performance. For convenience of the readers, we provide the details about this algorithm in this subsection.
@@
where $h(s_{uv})$ is the binary value over the weight $w_{uv}$ denoting whether this weight is selected and $s_{uv}$ the continuous score used in gradient descent based optimization. During backpropagation, the gradient will ignore $h(\cdot)$ and goes through it, therefore the gradient over $s_{uv}$ is
@@ A.3 Baselines
    Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017) applies a quadratic penalty over the parameters of a model based on their importance to maintain its performance on previous tasks.
@@
    Gradient Episodic Memory (GEM) (Lopez-Paz & Ranzato, 2017) selects and stores representative data in an episodic memory buffer. During training, GEM will modify the gradients calculated based on the current task with the gradient calculated based on the stored data to avoid updating the model into a direction that is detrimental to the performance on previous tasks.
@@
    Learning without Forgetting (LwF) (Li & Hoiem, 2017) is a knowledge distillation based method, which minimizes the discrepancy between the the old model output and the new model output to preserve the knowledge learned from the old tasks.
@@
    Bias Correction based Memory Replay (BCMR) (Chrysakis & Moens, 2023). This baseline is constructed by integrating the navie memory replay with the data sampling bias correction strategy  (Chrysakis & Moens, 2023). In other words, the method does not train the data immediately after observing the data. Instead, it stores the observed data into a memory buffer. Whenever testing is required, the model will be trained over all buffered data.
@@
    Scheduled Data Prior (SDP) (Koh et al., 2023) considers that the importance of new and old data is dependent on the specific characteristics of the given data, therefore balance the contribution of new and old data based on a data-driven approach. Since the condition assumed for $c^{2}$ in SDP (Koh et al., 2023) does not align with the actual situation in our experiments, the parameter $\alpha$ and $\beta$ is Eq. (9) of SDP) (Koh et al., 2023) are manually selected in our experiments.
@@ A.4 Model Evaluation
Different from standard learning setting with only one task to learn and evaluate, in our setting, the model will continually learn on a sequence of systems, therefore the setting and evaluation are significantly different. In the model training stage, the model is trained over a system sequence. In the testing stage, the model will be tested on all learned tasks. Therefore the model will have multiple performance corresponding to different tasks, and the most thorough evaluation metric is the performance matrix $\mathrm{M}^{p}\in\mathbb{R}^{N\times N}$, where $\mathrm{M}^{p}_{i,j}$ denotes the performance on the $j$-system after learning from the $1$-st to the $i$-th system, and $N$ is the number of systems in the sequence. In our experiments, each entry of $\mathrm{M}^{p}$ is a mean square error (MSE) value. To evaluate the overall performance on a sequence, the average performance (AP) over all learnt tasks after learning multiple tasks could be calculated. For example, i.e., $\frac{\sum_{j=1}^{i}\mathrm{M}^{p}_{N,j}}{N}$ corresponds to the average model performance after learning the entire sequence with $N$ tasks. Similarly, the average forgetting (AF) after $N$ tasks can be formulated as $\frac{\sum_{j=1}^{N-1}\mathrm{M}^{p}_{N,j}-\mathrm{M}^{p}_{j,j}}{N-1}$. These metrics are widely adopted in continual learning works (Chaudhry et al., 2018; Lopez-Paz & Ranzato, 2017; Liu et al., 2021; Zhang et al., 2023a; Zhou & Cao, 2021), although the names are different in different works.
For convenience, the performance matrix can be visualized as a color map (Figure 6). For example, they are named as Average Accuracy (ACC) and Backwarde Transfer (BWT) in (Chaudhry et al., 2018; Lopez-Paz & Ranzato, 2017), Average Performance (AP) and Average Forgetting (AF) in (Liu et al., 2021), Accuracy Mean (AM) and Forgetting Mean (FM) in (Zhang et al., 2023a), and performance mean (PM) and forgetting mean (FM) in (Zhou & Cao, 2021).